The length of a rectangle is four times its width.
If the area of the rectangle is 144cm squared , find its perimeter.

Answer by harpazo(655)

A = x(4x)
144 = 4x^2
144/4 = x^2
36 = x^2
Taking the square root of both sides yields
x = 6.
P = 2x + 2(4x)
Let x be 6 and solve for P to find your perimeter.
I hope this helps.